### v1

# Robot Drummer: Learning Rhythmic Skills for Humanoid Drumming

Abstract
Humanoid robots have seen remarkable advances in dexterity, balance, and locomotion, yet their role in expressive domains, such as music performance, remains largely unexplored. Musical tasks, like drumming, present unique challenges, including split-second timing, rapid contacts, and multi-limb coordination over pieces lasting minutes. In this paper, we introduce Robot Drummer, a humanoid system capable of expressive, high-precision drumming across a diverse repertoire of songs. We formulate humanoid drumming as sequential fulfillment of timed-contacts and transform drum scores in to a Rhythmic Contact Chain. To handle the long-horizon nature of musical performance, we decompose each piece into fixed-length segments and train a single policy across all segments in parallel using reinforcement learning. Through extensive experiments on over thirty popular rock, metal, and jazz tracks, our results demonstrate that Robot Drummer consistently achieves high F1 scores. The learned behaviors exhibit emergent human-like drumming strategies, such as cross-arm strikes, and adaptive sticks assignments, demonstrating the potential of reinforcement learning to bring humanoid robots into the domain of creative musical performance. Project page: robot-drummer.github.io
Index Terms:
Humanoid Control, Robotic Musicianship, Reinforcement Learning
I Introduction
Achieving human-like musical expressivity in robots presents a significant challenge, requiring not only robotic dexterity but also fine rhythmic control. Drumming, in particular, is a highly complex task that demands precise timing, high-frequency intermittent contacts, fast movements, and dynamic motor coordination. While robotic manipulation has seen impressive advancements in recent years, these are often confined to goal-driven tasks with well-defined end states [1]. In contrast, playing music is inherently process-driven, involving intricate timing, fluid motion, and the ability to adapt to changing rhythms. Success depends not on reaching a final configuration, but on sustaining an evolving sequence over a long temporal horizon. Even minor deviations in timing can profoundly affect the emotional quality and impact of a performance.
Reinforcement Learning (RL) has demonstrated remarkable success in enabling humanoid robots to master goal-driven tasks, such as getting up from the ground [2], standing up across diverse postures [3], and tracking agile motions [1, 4, 5]. In the domain of robotic musicianship, deep RL has been applied to piano playing with anthropomorphic dexterous hands [6], as well as to a simplified two-DoF underactuated drumming arm [7]. Despite recent efforts, the challenges posed by process-driven tasks like music remain largely unexplored by policy-learning methods. No prior work has investigated the application of RL to autonomously coordinate a humanoid’s body for drumming.
In this work, we aim to address this gap by introducing Robot Drummer, a system with humanoid robot that learns to play a variety of songs. The robots’ task is to perform a set of drumming patterns based on MIDI (Musical Instrument Digital Interface) transcription of sheet music, replicating the nuanced timing and expressivity typical of human drummers. This task requires our system to master three key challenges (i) temporal precision, ensuring each drum strike occurs at the exact moment required by the rhythm; (ii) spatial coordination, enabling the robot to strike multiple drums simultaneously or in quick succession; and (iii) dynamic adaptation, adjusting the motion in response to changes in tempo and musical intensity. While most of the current research in humanoid robotics has focused on loco-manipulation and whole-body control, this work aims to extend the frontier by endowing humanoid robots with creative and expressive skills.
II Related Works
II-A Learning for Humanoid Control
Recent humanoid controllers largely follow a motion imitation-based learning paradigm, where policies are trained to reproduce re-targeted reference trajectories. These are typically drawn from human motion-capture datasets [8] or human motion videos [9, 10, 5], often augmented by motion priors [11] and trained with curriculum schedules [3, 10] to enhance robustness and diversity. To tackle contact-rich and dynamic tasks such as parkour jumping, [12] introduced a sequential-contact decomposition approach that uses staged reference contacts to simplify reward design. Sim-to-real alignment efforts, such as those in [5], have improved fidelity for agile whole-body motion tracking policies in real environments. While these approaches have shown impressive performance in acquiring loco-manipulation skills, they are tied to demonstrated motions in the datasets, optimize primarily over a few seconds long-horizon, and lack explicit objectives for discovering novel, temporally precise contact sequences, which are essential for process-driven tasks such as drumming.
II-B Learning to Play Music
Robotic musicianship has advanced from purely trajectory-based systems [13] to data-driven and learning-based paradigms [6, 7]. Early works used simple arms and rules-based improvisation to interactively play alongside humans [14]. More recently, RL has enabled end‐to‐end training of anthropomorphic hands to play hundreds of piano pieces [6]. In drumming, a two-DoF underactuated arm was trained to generate rhythmic patterns under simplified kinematics [7]. Additionally, wearable robotic prosthetics have been developed to augment human drumming abilities, using muscle signals to adapt to the musical contexts [15]. Despite these advances, no prior work has investigated training a humanoid body to learn temporally precise, spatially coordinated drumming patterns.
To address this gap, we introduce Robot Drummer, a general RL framework for humanoid drumming that learns to perform drum patterns derived from MIDI transcriptions. In Robot Drummer, we formulate drumming as sequential fulfillment of timed contacts and decompose a full drum track into short, fixed‐length contact segments, trained in parallel with a single policy. We also design dense rhythmic contact rewards that treat each contact event (correct, wrong, or missed) as a learning signal, thereby guiding the policy toward correct rhythmic behaviors. Together, these components enable the training of policies that play drum tracks across a wide range of songs, bringing RL-driven humanoid control into the domain of expressive, process-driven musical performance.
Figure 2: Drum kit configuration used for training and evaluating the humanoid robot drummer (based on [16]). The setup includes a snare drum, two toms, a hi-hat, and two cymbals. Each drum is assigned a unique MIDI note, establishing a fixed note-to-drum mapping.
Figure 3: Overview of the Robot Drummer: Starting from a raw MIDI drum track (left), each note‐onset is first mapped to a specific drum and then converted into a sequence of timed ”contact events”, forming the Rhythmic Contact Chain (top center). Each contact in the chain is encoded as a one‐hot vector indicating spatiotemporal targets (which drum to hit at each timestep). This chain is then split into fixed‐length segments, each segment corresponding to a distinct portion of the song (bottom center) for parallel training. At each timestep, the agent’s state combines the next $L$ contact goals, the robot’s proprioception, and spatial context (positions of sticks and drums). The policy then outputs joint position targets for a proportional‐derivative controller (right).
III Robot Drummer: Learning Rhythmic Skills
Our main goal is to enable a humanoid robot to perform a wide range of songs by striking correct sequence of drums (as illustrated in Figure 2), based on MIDI transcription of sheet music. Each drum is modeled as a collision body which registers an impact when struck with a sufficient force. The synthesizer then produces a corresponding sound upon detecting drum strike. Drumming is a highly dynamic and contact-rich task that requires executing a series of precisely timed strikes over a long temporal horizon. The length of songs can range from few minutes to well over 6 minutes. The long horizon nature of the task combined with the variety of rhythmic patterns, and reward sparsity in high dimensional space makes RL exploration particularly challenging. In this work, we tackle these challenges by introducing two main approaches: (1) formulating drumming as a sequential fulfillment of contacts, which we refer to as Rhythmic Contact Chain; (2) temporal decomposition of a long-horizon musical piece into multiple segments, each trained in parallel with a unified policy. This formulation provides a general framework for learning rhythmic skills through task-agnostic contact rewards, improving sample efficiency of exploring in high-dimensional space. An overview of our approach is illustrated in Figure 3.
III-A Music Representation
While music can be represented as raw waveforms or spectrograms, these formats present challenges for robots, particularly when processing continuous data and designing appropriate reward functions. We use MIDI as a musical representation, which provides a standardized way to encode instruments’ timing in digital music. We begin with a MIDI file containing all the instruments of a musical piece. To isolate the percussion, we extract only the events on the dedicated drum channel, yielding a sequence of time-stamped “note-on” messages that correspond to drums hits. Each ”note-on” event contains pitch (MIDI note number), velocity, start time, and end time. From this drum-only track, we manually map each distinct MIDI note number to a physical drum in drum kit (see Figure 2). This mapping is song-specific and created after identifying the unique MIDI note pitches that are present in each song.
In MIDI, each drum in a drum kit has multiple articulations–variations of how the drum is played (e.g., open or closed hi-hat). These articulations produce distinct sounds and are represented as separate note numbers on a General MIDI percussion map  [^1] MIDI Percussion Map. When multiple articulations of the same drum appear in a given song, we select only the most frequent one for simplicity. While this approach may discard expressive variations for some songs, evaluation on the example tracks showed no significant drop in musical quality. Extending our framework to multiple articulations is left for future work.
After assigning the MIDI note numbers to each drum, we convert the raw MIDI data into a time-indexed drum sequence (or drum track), where each drum hit is encoded as a one-hot vector of length 6, indicating the drums that need to be struck at that timestep.
III-B Rhythmic Contact Chain
Humanoid drumming performance can be viewed as a sequence of precise and coordinated contact events, where each event corresponds to drum(s) strikes at a specific time. We model drumming as a chain of timed contact events, referred to as the Rhythmic Contact Chain (RCC), representing the temporal sequence of contact actions the robot must execute to perform a musical piece. Formally, we define it as:
$\mathcal{C}=\{\,C_{1},C_{2},\dots,C_{N}\},$ (1)
where $C_{i}$ denotes the contact step at time $t_{i}$ and may include multiple contact pairs corresponding to the target drum strikes. Each contact step is defined as:
$C_{i}=\bigl{(}D_{i},\,t_{i},\,S_{i}\bigr{)},$ (2)
with
$D_{i}\subseteq\mathcal{D},\quad S_{i}\subseteq\{s_{\text{L}},s_{\text{R}}\},$
where $D_{i}$ denotes the set of drums to strike at time $t_{i}$, as specified by the drum track (see Section III-A), $S_{i}$ specifies which stick(s)–left, right, can be used to execute the contacts, and $\mathcal{D}$ contains all the drums in a drum kit as shown in Figure 2. Each contact event in RCC thus corresponds to spatiotemporal targets derived from the drum track with flexible assignment of sticks. While RCC does not currently model strike velocity, it could easily be extended to include MIDI note velocities to capture dynamic intensity.
III-C Temporal Decomposition
Performing an entire drum track (several minutes long) in a single RL episode leads to sparse rewards and poor exploration, since thousands of timesteps must be rolled out sequentially before any reset, limiting parallelism and delaying exposure to later song sections. To address this, we introduce a temporal decomposition approach that partitions the full drum track into a sequence of fixed-length training segments, all trained in parallel. Each segment corresponds to a subset of contact steps in Rhythmic Contact Chain $\mathcal{C}$. Specifically, we divide $\mathcal{C}$ in to $M$ segments:
$\mathcal{C}^{(m)}\;=\;\bigl{\{}\,C_{i}:(m-1)\,P+1\leq i\leq m\,P\bigr{\}},$ (3)
where $P~{}=~{}N/M$ is the number of contact steps per segment, and $m=1,2,\dots,M$ represents the segments of a track. Each segment $\mathcal{C}^{(m)}$ defines a temporally localized portion of the song. This decomposition transforms a single long-horizon task into multiple shorter training segments, significantly improving the sample efficiency of learning rhythmic skills.
III-D Policy Training
We formulate humanoid drumming as a finite-horizon extended Markov Decision Process (MDP), defined by a tuple $(\mathcal{S,A,T,R,\gamma,C})$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{T}$ governs the transition dynamics, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{P}(\mathcal{S})$, $\mathcal{R}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $\gamma\in[0,1]$ represents the discount factor, and $\mathcal{C}$ is the contact chain defined in Section III-B. The objective is to learn an optimal policy $\pi_{\theta}$ that maximizes the expected cumulative reward $\mathbb{E}_{\pi_{\theta}}\left[\sum_{t=0}^{H-1}\gamma^{t}r_{t}\right]$ over the horizon $H$. To optimize the policy, we employ Proximal Policy Optimization (PPO) [17], an on-policy RL algorithm known for its stability and efficiency, particularly in large-scale parallel training environments.
III-D1 States
The state $\bm{s}_{t}$ is composed of three components: the robot’s proprioception $\bm{s}_{t}^{\mathrm{p}}$, spatial information related to the drums and sticks $\bm{s}_{t}^{\mathrm{sp}}$, and contact goals $\bm{s}_{t}^{\mathrm{c}}$. The proprioceptive state is defined as $\bm{s}_{t}^{\mathrm{p}}~{}\triangleq~{}[\,\bm{q}_{t},\;\dot{\bm{q}}_{t},\;\bm{a}_{t-1}\,],$ where $\bm{q}_{t}\in\mathbb{R}^{15}$ are the joint positions, $\dot{\bm{q}}_{t}\in\mathbb{R}^{15}$ are the joint velocities, and $\bm{a}_{t-1}\in\mathbb{R}^{15}$ is the action applied at the previous timestep. The spatial context $\bm{s}_{t}^{\mathrm{sp}}\in\mathbb{R}^{24}$ encodes the positions of the drums and sticks. The contact goals state is defined as $\bm{s}_{t}^{\mathrm{c}}~{}\triangleq~{}[\,\bm{d}_{t},\bm{d}_{t+1},\dots,\bm{d}_{t+L}\,],$ where $\bm{d}_{t+i}\in\mathbb{R}^{6}$ is a one-hot encoding of target drum strikes at timestep $t+i$ and $L$ is lookahead horizon over future steps to provide rhythmic context. The lookahead horizon controls how far into the RCC the robot can see. The contact goals state is akin to a time-phase variable in motion tracking policy learning [5] and effectively functions as a discrete phase marker, informing the policy ”which” drum(s) it should be targeting at that moment in the musical piece. It represents progression of the rhythm from start to end, serving as an implicit goal signal. We evaluate the impact of contact goals and the lookahead in our ablation study by comparing it to a continuous time phase variable.
III-D2 Actions
The policy’s actions $\bm{a}_{t}\in\mathbb{R}^{15}$ correspond to the target joint angles for a Proportional-Derivative (PD) controller, which applies torques to the robot’s joints. The torque at timestep $t$ is computed as:
$\bm{\tau}_{t}~{}=~{}\bm{K}_{p}(\bm{q}_{t}^{d}-\bm{q}_{t})-\bm{K}_{d}(\bm{\dot{q}_{t}}),$ (4)
where $\bm{q}_{t}^{d}$ is the desired joint configuration, defined as $\bm{q}_{t}^{d}~{}=~{}\beta\bm{a}_{t}+\bm{q}_{0}$; $\beta$ is an action scaling factor, $\bm{q}_{0}$ represents default joint positions of each joint, and $\bm{K}_{p}$ and $\bm{K}_{d}$ are the stiffness and damping coefficients of the PD controller, respectively.
III-D3 Rewards
Effective drumming requires executing hundreds of precisely timed strikes on the correct drums while avoiding any wrong hits. Optimizing solely with the sparse rewards fails to guide the policy toward right rhythmic behaviors. We hypothesize that every contact event–correct, wrong, or missed–provides crucial information for learning the fine‑grained timing and coordination needed in drumming. We thus propose unified task reward as:
$r_{t}=r_{t}^{\mathrm{contact}}+r_{t}^{\mathrm{reg}},$
where $r_{t}^{\mathrm{contact}}$ is the dense contact reward that encourages each correct contact while penalizing each wrong and missed contact, and $r_{t}^{\mathrm{reg}}$ encodes the regularization rewards. We define the contact reward as:
$r_{t}^{\mathrm{contact}}=w_{\mathrm{c}}\,r_{t}^{\mathrm{correct}}+w_{\mathrm{w}}\,r_{t}^{\mathrm{wrong}}+w_{\mathrm{m}}\,r_{t}^{\mathrm{missed}}+w_{\mathrm{p}}\,r_{t}^{\mathrm{prox}},$
where $r_{t}^{\mathrm{correct}}$ counts the correct strikes and rewards the policy for each drum impact that matches target drum(s), $r_{t}^{\mathrm{wrong}}$ penalizes the policy for hitting wrong drums at any time, $r_{t}^{\mathrm{missed}}$ penalizes the policy for failing to hit the target drums, and $r_{t}^{\mathrm{prox}}$ encourages the agent for moving sticks closer to target drums at moments when the hits are expected. The first three terms directly correspond to the contact events in Rhythmic Contact Chain, defined in Section III-B. A summary of all reward terms with their respective weights is provided in Table I. The weights are based on an intuition that missing a strike is worse than hitting a wrong drum, while correct hits should be rewarded more to encourage well-timed strikes.
TABLE I: Reward Components
Term Expression Weight
Contact
Correct strikes $\sum_{e\in\mathcal{E}_{t}}\mathbb{I}\{\text{drum}(e)\in D_{i}\}$ $+1.0$
Wrong strikes $\sum_{e\in\mathcal{E}_{t}}\mathbb{I}\{\text{drum}(e)\notin D_{i}\}$ $-0.5$
Missed strikes $\sum_{d\in D_{i}}\mathbb{I}\{\text{drum missed}=d\}$ $-2.0$
Proximity $\mathbb{I}\{|D_{i}|>0\}\sum_{s\in\{s_{\text{L}},s_{\text{R}}\}}min_{d\in D_{i}}\|p_{s,t}-p_{d}\|$ $-1.0$
Regularization
Action‑rate $\|a_{t}-a_{t-1}\|^{2}$ $-1e^{-3}$
DoF acceleration $\|\ddot{q}\|^{2}$ $-2.5e^{-7}$
  • 1
    Note: $\mathcal{E}_{t}$ is the set of drums the robot actually hits at time $t$; $p_{s,t}$ is the position of stick $s$ at time $t$; $p_{d}$ is the position of drum $d$.
Figure 4: Sequential snapshots of a complex drum-fill segment from ”Linkin Park – In the End”, showing the humanoid robot executing multiple strikes on four drums in under two seconds. From left to right: the robot raises both sticks, strikes the hi-hat precisely, executes a simultaneous hi-hat and snare hit, transitions to a tom 2 hit, and finishes with a cymbal 1 strike. During the 2 seconds interval, the robot makes multiple rapid contacts on the drums, but these frames show only the representative moments of the fill.
III-E Music State Initialization
State initialization plays a crucial role in training RL policies, where the key idea is to initialize the agent’s starting state at random phases along a reference motion trajectory. Recent works have shown the effectiveness of state initialization in tracking reference motions [5, 18]. To apply this idea in the context of drumming, we propose Music State Initialization, where we randomly sample contact steps within a segment $\mathcal{C}^{(m)}$ of RCC and initialize the corresponding contact goals state at those steps. This effectively randomizes the starting point of the drum track and exposes the policy to high reward contact states early on in the training.
IV Experiments
In this section, we present comprehensive experiments designed to evaluate the performance of the Robot Drummer. We conduct our experiments on a Unitree G1 humanoid robot in Isaac Gym [19] simulation environment, where we control 15 degrees of freedom (DOF) in robot’s upper body–7 DOF for each arm and 1 DOF for torso. Our experiments are designed to address the following key questions:
  • Q1. How does the Robot Drummer perform across songs of varying complexity?
  • Q2. What role does each component of the Robot Drummer’s design play in its overall performance?
  • Q3. Is the Robot Drummer capable of learning a single generalist policy that enables it to play multiple songs?
IV-A Quantifying Rhythmic and Spatial Complexity
To address Q1, we first identify the factors that affect the difficulty of drumming in a given song and then train specialist policies. The primary difference in songs complexity lies in the pattern of drum hits, both in time and space. To systematically assess these patterns, we introduce a set of features that capture the song’s key characteristics, such as beats per minute (BPM), normalized Pairwise Variability Index (nPVI) of the inter-onset intervals (IOI), time signature changes, the number of distinct drums used, and hit distribution entropy. These features are described in detail below and collectively provide a quantitative measure of the song’s complexity.
TABLE II: Quantitative evaluation of policies on songs with zero polyphony
Song #Drums Entropy nPVI BPM F1 $\uparrow$
David Bowie - Rebel Rebel 2 0.27 8.76 130 0.985 $\pm$ 0.008
The White Stripes - Seven Nation Army 4 0.97 6.71 120 0.977 $\pm$ 0.004
Nirvana - Lithium 6 0.74 22.10 112 0.954 $\pm$ 0.018
Jimi Hendrix - Fire 6 0.67 29.48 156 0.943 $\pm$ 0.021
Dave Brubeck - Take Five 4 0.71 51.86 167 0.908 $\pm$ 0.034
Linkin Park - In The End 6 0.61 32.68 86 0.901 $\pm$ 0.020
Bon Jovi – Livin’ on a Prayer 6 0.84 82.93 123 0.885 $\pm$ 0.029
The Police - Roxanne 6 0.73 21.19 136 0.878 $\pm$ 0.024
  • 1
    All songs have zero time signature changes, except for ”Bon Jovi – Livin’ on a Prayer” (4).
IV-A1 Beats per Minute
The Beats per minute (BPM) measures the tempo or speed of the song. It indicates how fast or slow the song is played, determining the pace at which drumming occurs.
IV-A2 Number of Drums Used
The number of distinct drums in a song directly affects its rhythmic and spatial complexity. More drums introduce variation into the pattern, requiring greater coordination and contributing to a more intricate rhythm. In this work, the maximum number of drums that can be used in a song is 6, as illustrated in Figure 2.
IV-A3 Time Signature Changes
Time signature changes refer to shifts in the song’s rhythmic structure, such as changes in how many beats are present in each measure. The changes in time signature affect the complexity of the song, as it requires the robot to adapt to different rhythmic patterns throughout the piece.
IV-A4 Normalized Pairwise Variability Index (nPVI)
The normalized Pairwise Variability Index (nPVI) quantifies the relative variability in timing between successive drum hits, known as inter-onset intervals (IOIs). For each adjacent pair of IOIs, we compute the absolute difference divided by their mean, then average these values across the entire song. A higher nPVI reflects greater temporal irregularity and thus a more challenging rhythm for the robot to execute. This metric is widely used in music analysis to capture rhythmic complexity across diverse musical styles [20].
IV-A5 Hit Distribution Entropy
Some songs concentrate drum hits on a few drums, while others spread them more evenly. We use Shannon entropy to quantify how dispersed the strikes are. A more even distribution of drum hits yields higher entropy, indicating that the robot must switch between different drums throughout the song and coordinate wider spatial movements between the hits.
IV-A6 Polyphony
Although our humanoid robot is physically limited to striking two drums at once (one per stick), MIDI drum sequences often encode complex polyphonic textures with three or more drums at the exact same moment, for example, a simultaneous hit on the snare, tom, and cymbal. We quantify this polyphony by measuring the fraction of time‐steps where the MIDI drum track requests more than two simultaneous hits. These overlapping notes signal moments in the song when the robot must prioritize which drums to hit.
IV-B Evaluation Metrics
To evaluate the drumming performance of learned policies, we use precision, recall, and F1 score by comparing ground‐truth sequence of desired drum strikes in RCC (see Section III-B) to the robot’s actual hits at each timestep. Precision measures the proportion of the robot’s strikes that correctly align with RCC events (i.e., on the intended drums) over the entire song, while recall captures the fraction of ground-truth contact events the robot actually executes. The F1 score is the harmonic mean of precision and recall and provides a single, balanced metric of overall rhythmic fidelity on a 0-1 scale. We primarily base our evaluations on F1 score.
Note that the presence of polyphony in a song will inevitably reduce the overall scores, as the robot is physically unable to execute all the simultaneous contact events in the RCC.
IV-C Policy Training and Evaluation Protocol
We use a 3-layer MLP architecture for both the actor and critic networks, with hidden sizes of (128, 64, 32). For each song or training condition, policies are learned across 5 random seeds over 1500 PPO updates. Experiments are conducted on a server with Nvidia RTX 4090 GPUs, running 8192 parallel environments in Isaac Gym. After training, each policy is evaluated over 20 independent rollouts, and performance is measured using the mean F1 score between actual and target drum strikes across all timesteps. The final performance metric is the average F1 score and standard deviation across the 5 seeds.
Figure 5: (a) Relative influence of song metrics on drumming performance across all specialist policies. Bars show the absolute Spearman rank correlation between each song‐level feature and F1 score. Rhythmic variability (nPVI), Number of drums, and polyphony have the largest correlations, with polyphony imposing an upper‐bound limit on performance (robot can hit at most two drums simultaneously). Entropy and tempo exhibit comparatively weaker standalone effects. (b)–(d) Scatter plots with linear trendlines illustrating how increasing nPVI (b) and drum count (c) correspond to lower F1 scores, while entropy (d) shows a slight downward trend, indicating that rhythmic diversity has little impact on its own.
Refer to captionRefer to captionRefer to caption
Figure 6: Sequence of frames, illustrating a robot performing a fast drum roll (within a one-second interval) on the snare drum during the songs ”Nirvana - Come As You Are” (top figure) and ”Foo Fighters - Everlong” (bottom figure). The frames highlight how the policies precisely position the right hand to execute rapid strikes.
Figure 7: Performance comparison of Nirvana - Lithium (top) and The Police - Roxanne (bottom) policies during an RL episode, showing how target drum hits align with executed strikes (left) alongside the original drum track (right). The plot shows that Lithium’s regular, well‐spaced drum events allow the robot to hit nearly every target, whereas Roxanne’s irregular, densely interleaved hits force rapid spatial switches that the policy cannot execute, leading to missed cymbal 1 and tom 2 strikes and thus a lower F1 score. The drum track is plotted at 10 fps for visualization, while the policies use a 50 fps discretization.
TABLE III: Ablation of Reward and Observation Components. F1 scores are reported as mean $\pm$ std over 5 runs.
Reward / Observation Components Song
Reward Ablation
Correct Strikes Wrong Strikes Missed Strikes Rebel Rebel Lithium Fire In The End Livin’ on a Prayer Roxanne
0.64 $\pm$ 0.16 0.70 $\pm$ 0.04 0.37 $\pm$ 0.06 0.60 $\pm$ 0.05 0.35 $\pm$ 0.04 0.52 $\pm$ 0.04
0.96 $\pm$ 0.00 0.86 $\pm$ 0.03 0.91 $\pm$ 0.01 0.70 $\pm$ 0.04 0.62 $\pm$ 0.16 0.69 $\pm$ 0.00
0.99 $\pm$ 0.00 0.95 $\pm$ 0.02 0.94 $\pm$ 0.02 0.90 $\pm$ 0.02 0.89 $\pm$ 0.03 0.88 $\pm$ 0.02
Observation Ablation
Proprioception & Spatial Phase Variable Contact Goals Rebel Rebel Lithium Fire In The End Livin’ on a Prayer Roxanne
0.07 $\pm$ 0.03 0.07 $\pm$ 0.02 0.03 $\pm$ 0.00 0.06 $\pm$ 0.01 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00
0.12 $\pm$ 0.02 0.09 $\pm$ 0.03 0.05 $\pm$ 0.00 0.10 $\pm$ 0.00 0.01 $\pm$ 0.00 0.02 $\pm$ 0.00
0.99 $\pm$ 0.00 0.95 $\pm$ 0.02 0.94 $\pm$ 0.02 0.90 $\pm$ 0.02 0.89 $\pm$ 0.03 0.88 $\pm$ 0.02
TABLE IV: Quantitative evaluation of policies on songs with non-zero polyphony
Song #Drums Entropy nPVI BPM Polyphony (%) F1 $\uparrow$
Limp Bizkit - Break Stuff 4 0.83 10.22 107 0.27 0.977 $\pm$ 0.005
Metallica - Nothing Else Matters 5 0.61 5.08 69 0.23 0.962 $\pm$ 0.003
Foo Fighters - Everlong 4 0.77 3.79 156 0.86 0.955 $\pm$ 0.017
Metallica - One 5 0.64 11.74 142 1.21 0.953 $\pm$ 0.006
AC/DC - You Shook Me All Night Long 6 0.45 2.05 129 1.71 0.945 $\pm$ 0.017
Europe - The Final Countdown 6 0.60 48.96 118 1.56 0.940 $\pm$ 0.011
Rage Against the Machine - Killing in the Name 6 0.67 24.38 102 0.14 0.940 $\pm$ 0.006
Eagles - Hotel California 6 0.54 5.38 149 0.98 0.933 $\pm$ 0.020
Iron Maiden - Where Eagles Dare 5 0.64 34.82 120 1.16 0.923 $\pm$ 0.016
Green Day - Boulevard of Broken Dreams 6 0.76 7.96 167 0.20 0.920 $\pm$ 0.009
Nirvana - The Man Who Sold the World 5 0.57 17.08 118 0.40 0.919 $\pm$ 0.017
Nirvana - Come as You Are 6 0.57 8.73 120 0.71 0.916 $\pm$ 0.013
Nirvana - Dumb 6 0.90 14.59 120 1.60 0.915 $\pm$ 0.050
Metallica - Enter Sandman 5 0.86 10.63 123 0.55 0.908 $\pm$ 0.025
The Police - Message in a Bottle 6 0.64 14.55 150 0.21 0.906 $\pm$ 0.021
Guns N’ Roses - Sweet Child o’ Mine 5 0.67 13.66 128 0.21 0.902 $\pm$ 0.040
AC/DC - Thunderstruck 6 0.45 27.17 107 3.36 0.895 $\pm$ 0.013
Guns N’ Roses - Knockin’ on Heaven’s Door 6 0.73 79.42 128 0.31 0.880 $\pm$ 0.023
Iron Maiden - Wasted Years 5 0.58 2.02 152 4.58 0.863 $\pm$ 0.013
Bon Jovi - Bed of Roses 6 0.74 36.49 60 0.64 0.849 $\pm$ 0.011
Bon Jovi - It’s My Life 6 0.56 92.40 120 1.69 0.825 $\pm$ 0.022
Bon Jovi - Always 6 0.73 23.61 67 3.28 0.814 $\pm$ 0.020
Rolling Stones - Paint It Black 5 0.97 91.76 110 3.20 0.792 $\pm$ 0.055
Nirvana - Smells Like Teen Spirit 6 0.77 22.02 119 14.16 0.712 $\pm$ 0.030
  • 1
    All songs have zero time signature changes, except for ”Metallica - Nothing Else Matters” (17), and ”Nirvana - The Man Who Sold the World” (4).
V Results
V-A Specialist Polices
To evaluate the performance of Robot Drummer, we trained over 30 specialist policies on popular rock and metal tracks of varying complexities. Table II and  IV present the complexity metrics for songs with zero polyphony (no more than 2 simultaneous hits) and non‐zero polyphony, respectively, alongside corresponding F1 scores achieved by the policies. The results show that specialist policies consistently achieve high F1 scores. Fewer drums and regular rhythmic patterns (low nPVI) yield near-perfect performance, while increased drum count and higher rhythmic variability (high nPVI) drops the F1 score. Additionally, Figures 4 and  6 provide qualitative snapshots of challenging drumming sequences: a drum fill, which involves a rapid succession of hits across multiple drums to transition between sections, and a snare roll, which is a high-speed hits on a snare drum. These examples demonstrate the robot’s ability to perform both spatially and temporally demanding patterns with precise coordination. Full videos of drumming performances are available on our project website.
To quantify which song characteristics most strongly affect performance, we also computed Spearman’s rank correlation between each feature and per-song F1 score across all specialist policies in Figure 5. Among all factors, rhythmic irregularity (nPVI) shows the strongest negative correlation with F1 score, indicating that songs with highly uneven timing between hits are most challenging. The number of drums used in a song also correlates negatively with F1 score. This trend aligns with our expectations as each additional drum target amplifies the difficulty of policy learning due to a larger spatial manifold. The policy needs to coordinate larger movements across a wider space with a greater potential for incorrect and missed hits. Polyphony, on the other hand imposes a fundamental cap on upper-bound performance, since the robot is physically limited to executing at most two strikes simultaneously. Thus, songs with high polyphony (i.e., frequent three or four-drums simultaneous contacts) inevitably incur missed hits. Interestingly, entropy (a measure of rhythmic diversity) showed minimal impact on performance. This suggests that the policies can reliably handle more varied drum usage in a song as long as the timing is consistent. Similarly, tempo did not significantly affect outcomes; fast songs were not inherently harder, provided the rhythm was regular. Finally, we observed negligible correlation between time signature changes and F1 score, due to the fact that only three songs involved the changes in time signature.
Figure 7 compares the performance of specialist policies on two representative songs (Nirvana - Lithium and The Police - Roxanne), by overlaying the target hits versus actual robot strikes alongside the original drum track. The two policies achieve notably different F1 scores despite similar values for all rhythmic complexity metrics. Lithium’s drum pattern is highly regular and evenly spaced, whereas Roxanne’s pattern has irregular cluster of hits with many switches between drums in quick succession (see Figure 7 (right)). The regularity in Lithium gives the robot sufficient time to swing each stick between target drums, resulting in an almost perfect F1 score (0.977). By contrast, Roxanne’s sequence interleaves distinct drum targets in a short time windows. For example, the transition from cymbal 2 to tom 2 occurs rapidly, and as a result, the policy stays near cymbal 2 and never attempts to hit tom 2. Likewise, cymbal 1 hits are squeezed into very narrow gaps between hi-hat strikes, and the policy tends to prioritize more frequent pattern, staying locked in on hi-hat while missing the majority of cymbal 1 strikes. These examples reflect how the policies learn a preference for spatially efficient actions under tight timing constraints.
V-B Ablation Analysis
To assess the individual contributions of our design choices, we perform an ablation study over the contact reward components and the observation modalities, reporting mean and standard deviation of F1 scores across five independent runs in Table III. In the reward ablation, we retain proximity and regularization terms and compare three variants: (i) correct-strike reward only, (ii) correct plus wrong strikes, and (iii) the full dense contact reward (correct + wrong + missed) strikes. The results demonstrate that the full reward formulation is essential for achieving robust rhythmic accuracy, as omitting either wrong or missed strike terms substantially degrades F1 across all five songs.
We next evaluate the effect of removing key components from the policy’s input state. We find that the policies trained without the one-hot contact goals collapse to near-zero F1, indicating they lack the ability to anticipate upcoming drum strikes. We also examine the role of an explicit time phase signal by replacing the one-hot contact goals with a continuous, normalized phase variable [18] ranging between $0-1$. Phase-conditioned policies likewise perform poorly, suggesting that coarse temporal context alone is insufficient to acquire drumming skills. The time phase variable informs the policy ”how far through the segment” it is but not ”which drum is next”, highlighting the necessity of explicit goal encoding. Collectively, these results show that humanoid drumming requires both dense contact rewards and structured state representation that encode future rhythmic targets.
Finally, we evaluate the impact of removing our temporal decomposition scheme by training policies end-to-end on full-length tracks rather than on fixed-length segments. Results are summarized in Table V. While the final F1 scores remain comparable in both settings, we observe a substantial increase in wall-clock training time: policies trained without decomposition require approximately 8–9 hours to converge, compared to just 2–3 hours with temporal segmentation on the same hardware.
TABLE V: F1 Scores With and Without Temporal Decomposition
Song w/o Decomposition w/ Decomposition
Rebel Rebel 0.98 $\pm$ 0.01 0.99 $\pm$ 0.00
Lithium 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02
Fire 0.93 $\pm$ 0.02 0.94 $\pm$ 0.02
In The End 0.92 $\pm$ 0.02 0.90 $\pm$ 0.02
Livin’ on a Prayer 0.86 $\pm$ 0.02 0.89 $\pm$ 0.03
Roxanne 0.89 $\pm$ 0.02 0.88 $\pm$ 0.02
V-C Generalist Policies
We also trained generalist policies to play multiple songs with a single controller. Specifically, we selected six songs in Table II (except Seven Nation Army, and Take Five) and trained a single policy on them. As expected, the generalist policy performed markedly worse on each song than specialist policies as shown in Figure 8. Songs, such as, Bon Jovi – Livin’ on a Prayer and Jimi Hendrix - Fire experienced the largest drop in F1 scores, while more regular or slower songs like Rebel Rebel (0.88) and In The End (0.89) were comparatively less affected. By contrast, specialist policies on these songs achieved F1 values near or above 0.9. Training on all eight songs in Table II further degraded performance on every track, indicating that naive multi-song learning forces the policy into suboptimal compromises.
Combining multiple rhythmic patterns (e.g. irregular timing of Livin’ on a Prayer and Fire’s fast tempo) in one policy introduces negative transfer, impacting the policy’s ability to generalize. Several factors likely contribute to this generalization difficulty. The songs vary greatly in their spatial layout (e.g. 2-6 drums) and temporal structure, so the policy must accommodate conflicting demands. As our specialist results show, increasing the number of drums and temporal irregularity already correlate with lower performance. The generalist policy must cover both simple and complex patterns, which appears to impact it’s ability to perform well on a specific pattern. We observed that the generalist policies tend to adopt ’safer’ strategies, for instance, they often focus on repeatedly used drum targets and ignores others, much like the Roxanne’s specialist policy in Section V-A.
V-D Listener Study
To evaluate the artistic and expressive qualities of Robot Drummer, we conducted a small-scale listener study. Fifteen participants each watched and listened to at least three of the six performance videos (80 seconds each, including ”Nirvana - Lithium,” ”Bon Jovi - Livin’ on a Prayer,” ”Eagles - Hotel California,” ”Metallica - One,” ”Linkin Park - In the End,” and ”The Police - Roxanne”) of specialist policies. After watching the videos, participants answered a series of questions, rating the robot’s drumming on a 1–5 Likert scale (1 = not at all; 5 = very much) for the following aspects:
  • Timing Consistency - How consistent and well-timed was the rhythm of the drumming?
  • Expressiveness - To what extent did the robot convey a sense of musical feel or expressiveness?
  • Naturalness - How natural or human‑like did the robot’s drumming sound?
  • Enjoyment - How enjoyable were the drumming performances overall?
Across all responses, we observed the following results:
  • Timing Consistency: $3.69~{}\pm~{}0.63$
  • Expressiveness: $3.38~{}\pm~{}0.96$
  • Naturalness: $2.46~{}\pm~{}0.78$
  • Enjoyment: $3.08~{}\pm~{}1.60$
These results show that listeners generally perceived the robot’s timing and musical feel in the upper mid-range, with lower ratings for human-like naturalness and mixed enjoyment responses reflected by greater variance. This brief study provides qualitative insight into how audiences experience the songs performed by Robot Drummer.
Figure 8: F1 scores for specialist versus generalist policies on selected songs, illustrating the performance drop when a single policy is trained across multiple rhythms.
VI Discussion
Our experiments highlighted both the strengths and limitations of RL in expressive humanoid drumming. Specialist policies mastered songs with high rhythmic fidelity, often with F1 scores above 0.9. The robot not only kept perfect timing during drumming performances but also executed complex drums fills and rolls (see Figures 4 and  6). Notably, the learned behaviors showed emergent spatial strategies: for example, the robot performed multiple cross-over strikes (hitting cymbal with the right stick and the left hand reaching for snare) while performing Livin’ on a Prayer. Such maneuvers were not explicitly encoded in the reward; they arose as the policy searched for efficient ways to reach distant drums.
Our analysis shows that spatial complexity (number of drums) and temporal variability (nPVI) are the dominant factors limiting performance. Songs with many quick switches between drums caused the robot to favor the most frequent targets and miss others, whereas metrics like drum‑hit entropy or tempo had much smaller effects. Moreover, training a single policy across multiple songs led to a drop in F1 scores, highlighting the inherent challenges of multi‑task RL, and suggesting that blending sequences with different rhythmic structures may require richer models or better training schemes. These findings imply that success depends critically on aligning arm trajectories with precise hit timings and requires planning in to the future (how the robot hits one drum directly influences its ability to reach the next targets). Our dense contact rewards and future goals encoding implicitly accounted for this. Indeed, we observed policies dynamically reassign sticks based on the musical context. For instance, in Livin’ on a Prayer, the policy typically used the right stick for the snare (a spatially efficient choice) but switched to the left stick when a cymbal 2 hit immediately followed the snare hit, ensuring the right hand remained available for the cymbal while still executing the snare strike on time. Without explicit future goal signals and contact rewards, the robot cannot anticipate upcoming hits and fails to coordinate its sticks effectively. Thus, rhythmic skill learning requires coupling of fine temporal context with spatial planning - akin to how a human drummer reads a music sheet or feels the beat and plans in to the future.
VI-A Pathway to Sim-Real Transfer
To effectively drum on the real hardware, the policies first need to be trained with sim-real techniques [5] such as domain randomization, random observation/action noise, and use of state history, to better handle real-world variability. Second, the hand-stick interface needs to be realized with the elastic material to have compliant dynamics. The elastic material provides just enough compliance to emulate human grip relaxation while keeping the stick firmly in place. During transfer, one potential path might involve humanoid playing on electronic drum-kit with velocity-sensitive pads (and optionally positional-sensing). The kit will output MIDI note-on events with velocity values for each hit that can be compared to the target MIDI sequence, yielding timing and force reward that fine‑tunes the policy and compensates for residual sim‑real mismatch. Optionally, we could also feed target velocity for each strike into the agent’s observations or directly into the reward to encourage behaviors that will produce more accurate sounds right from the beginning. Together, these measures could provide a practical, lightweight sim‑to‑real strategy tailored for drumming without extensive system identification.
VII Conclusion
In this work, we have introduced Robot Drummer, a general RL-based system that enables a humanoid robot to perform complex drumming patterns. Our key contributions include modeling drumming as a rhythmic contact chain with a temporal segmentation approach and dense contact rewards that together allow training on long rhythmic sequences. The learned policies achieve near-perfect timing on many songs and exhibit human-like strategies (e.g., cross-overs strikes, fast drum rolls, and dynamic sticks assignments) without explicitly encoding them. Through quantitative experiments, we identified key factors affecting performance: songs with many drums or highly irregular timing are substantially harder to master. We also explored the possibility of learning a generalist drumming policy on multiple songs. While many specialist policies achieved near perfect performance on individual songs, a single generalist policy scored much lower, highlighting the challenge for generalization in rhythmic control. Despite the challenges for multi-task RL, the specialist policies demonstrated remarkably fluid and expressive drumming skills in simulation, suggesting that RL-based humanoid control can bridge into the creative domain. In future work, we will bridge the sim‑to‑real gap by deploying Robot Drummer on physical hardware and investigate improvisation and style adaptation, enabling new forms of musical collaboration between robots and human musicians.
VIII Acknowledgments
This work benefited from the insightful discussions with Mirko Nava, and from the assistance of Marco Maccarini in setting up the necessary resources for training infrastructure that supported this research, both at IDSIA.